Find the equation (in standard form) of the line through (1, –7) and (4, –1). Match your result to the correct answer below.
A) 2x + y = 9
B) x-2y=18
C) 2x – y = 9
D) x-2y=9

Answer by Alan3354(69443)

Find the equation (in standard form) of the line through (1, –7) and (4, –1). Match your result to the correct answer below.
A) 2x + y = 9
B) x-2y=18
C) 2x – y = 9
D) x-2y=9
------------
Determinants is the easiest way
|x y 1|
|1 -7 1| = 0
|4 -1 1|
------
x*(-7 - (-1)) - y*(1-4) + 1*(-1 - (-28)) = 0
-6x + 3y + 27 = 0
2x-y = 9 (C)